# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

### v2

@@ B.3 Data Format
Taking the first piece as example , we express our data in json format with six keys
@@
"original_prompt": "a live recording of ambient acoustic
@@
"Editing_prompt" is the edited caption while "Original_prompt" is the origin one . "Blended_word" is the subject to be edited."Emphasize" is the word which should be emphasized. "editing_instruction" is the description of editing.Something else that needs attention is in the editing type "delete instrument" , we create another key "neg_prompt" , which helps to reduce the possibility of deleted instruments still appearing.
@@ Appendix C Implementation Details
For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (Liu et al., 2023b). Our assessment employs a comprehensive set of metrics, namely CLAP, LPAPS, Structure Distance, and FAD. These metrics are calculated using the CLAP models available in the AudioLDM_eval package, which is accessible at https://github.com/haoheliu/audioldm_eval. In line with the methodology described by  Manor & Michaeli (2024), we apply a forward guidance of 3 and a reverse guidance scale of 12 for DDPM inversion. For the DDIM inversion, the guidance scale is set to 5, while for SDEdit, we employ a guidance scale of 12. The forward guidance of DIC is 1 while the reverse scale is 5. We chose these values by exploration about different guidance scale, as discussed in Appendix D. We conduct all experiments in NVIDIA 4090. For our evaluation, we have selected the public pre-trained AudioLDM 2-Music model (Liu et al., 2023b). To ensure a thorough and multidimensional assessment, we measure performance using a suite of metrics that includes CLAP, LPAPS, Structure Distance, and FAD, and conduct on an NVIDIA 4090. The computation of these metrics is facilitated by the CLAP models provided within the AudioLDM_eval package, which is publicly available at https://github.com/haoheliu/audioldm_eval.
@@ C.1 Metrics
There are details about five metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) Struture Distance (Ju et al., 2023): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio. (2) CLAP Score (Elizalde et al., ): This criterion evaluates the degree to which the output conforms to the specified target prompt. (3) LPAPS (Iashin & Rahtu, 2021; Paissan et al., 2023): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), this measure evaluates the consistency of the edited audio with the original audio. (4) MSE (Mean Squared Error): This distance metric quantifies the variance between the source and the edited audio. (5) FAD (Fréchet Audio Distance) (Kilgour et al., 2018): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals.
@@ Appendix D Quantitative Results
For the lack of systematic experiments which determines the optimal combination of guidance scales for achieving the best editing performance, and analysis on how these guidance scales affect final consequence in both reconstruction and editing, we conduct this experiment to find the best scales.
@@ Appendix F Limitations
In spite of remarkable outcome of our method , but due to the limitation of the generation model we used , we are incapable of instigating a profound change.
Though our method is faster than est-optimization and end-to-end diffusion models,but due to the numerous steps it requires (T=200), the duration of computing distance is quite long.Thus , we will implement a more powerful text-to-music generation model to support better editing, while trying to use consistency model or flow matching model to achieve high quality and fast music generation in future work
@@ G.1 Change Instrument
In the following figures , we show the capability of MEDIC to change the instrument.Here we edit the ground truth music piece with original prompt "a live recording of ambient acoustic [guitar] music" and editing prompt "a live recording of ambient acoustic [violin] music". The difference of instruments can be observed in the Mel-spectrum
@@ G.2 Add Instrument
In the following figures , we show the capability of MEDIC to add more instrument.Here we edit the ground truth music piece with original prompt "a heavy metal instructional audio with a distortion guitar" and editing prompt "a heavy metal instructional audio with a distortion guitar [and drums]". The appearance of new instrument can be observed in the Mel-spectrum which present drum sound of high frequency.
@@
Figure 6: Editing Type 1 :Add Instrument
@@ G.3 Delete Instrument
In the following figures , we show the capability of MEDIC to delete instrument.Here we edit the ground truth music piece with original prompt "a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood" and editing prompt "a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood". The vanishing of instrument can be observed in the Mel-spectrum
@@
Figure 7: Editing Type 2 :Delete Instrument
@@ G.4 Change Genre
In the following figures , we show the capability of MEDIC to change genre of a music piece. Here we edit the ground truth music piece with original prompt "a recording of a solo electric guitar playing [blues] licks" and editing prompt "a recording of a solo electric guitar playing [rock] licks". The obvious difference of genre can be observed in the Mel-spectrum
@@
Figure 8: Editing Type 3 :Change Genre
@@ G.5 Change Mood
Mood is an important attribute of a music. In the following figures ,we show the capability of MEDIC to change mood of a music piece. Here we edit the ground truth music piece with original prompt "a recording of [aggressive] electronic and video game music with synthesizer arrangements"and editing prompt "a recording of [peaceful] electronic and video game music with synthesizer arrangements". The change of Mood can be observed in the Mel-spectrum
@@
Figure 9: Editing Type 4 :Change Mood
@@ G.6 Change Rhythm
Rhythm represents the speed of the music. In the following figures ,we show the capability of MEDIC to change Rhythm of a music piece. Here we edit the ground truth music piece with original prompt "a [slow] tempo ukelele tuning recording with static"and editing prompt "a [fast] tempo ukelele tuning recording with static". The change of Rhythm can be observed in the Mel-spectrum.The edited Mel-spectrum is much intensive.
@@
Figure 10: Editing Type 5 :Change Rhythm
@@ G.7 Change Background
In the following figures ,we show the capability of MEDIC to change background of instrument of a music piece. Here we edit the ground truth music piece with original prompt "an amateur ukulele recording with a [medium to uptempo] pace"and editing prompt "an amateur ukulele recording with a [steady and rhythmic] pace". The change of instrument background can be observed in the Mel-spectrum.
@@
Figure 11: Editing Type 6 :Change Background
@@ G.8 Change Melody
In the following figures ,we show the capability of MEDIC to change melody of a music piece. Here we edit the ground truth music piece with original prompt "this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo"and editing prompt "this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo". The change of Melody can be observed in the Mel-spectrum
@@
Figure 12: Editing Type 7 :Change Melody
@@ G.9 Extract Instrument
In the following figures ,we show the capability of MEDIC to extract one certain instrument of a music piece. Here we edit the ground truth music piece with original prompt "a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]"and editing prompt "a reggae rhythm recording with bongos". The change of instruments can be observed in the Mel-spectrum. Much sounds vanished.
@@
Figure 13: Editing Type 8 :Extract Instrument
@@
In the processing of the data and models involved in this study, we fully considered the potential risks. We ensure that all data sources are rigorously screened and vetted,and the model we used is absolutely trained from the safe dataset to minimize security risks of being misused.

### v3

@@ B.3 Data Format
Taking the first piece as an example, we express our data in JSON format with six keys
@@
"source_prompt": "a live recording of ambient acoustic
@@
“Editing_prompt" refers to the edited caption, while “source_prompt" denotes the original caption. "Blended_word" indicates the subject to be edited, and “Emphasize" represents the word that should be highlighted. “Editing_instruction" provides a description of the editing process. Additionally, in the editing type “delete instrument," we introduce another key, “neg_prompt", which helps reduce the likelihood of deleted instruments reappearing.
@@ Appendix C Implementation Details
For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (Liu et al., 2023c). Our assessment employs a comprehensive set of metrics, namely CLAP, LPAPS, Structure Distance, and FAD. These metrics are calculated using the CLAP models available in the AudioLDM_eval package, which is accessible at https://github.com/haoheliu/audioldm_eval. In line with the methodology described by  Manor & Michaeli (2024), we apply a forward guidance of 3 and a reverse guidance scale of 12 for DDPM inversion. For the DDIM inversion, the guidance scale is set to 5, while for SDEdit, we employ a guidance scale of 12. The forward guidance of MEDIC is 1 while the reverse scale is 5. We chose these values by exploring different guidance scales, as discussed in Appendix D. We conduct all experiments in NVIDIA 4090. For our evaluation, we have selected the public pre-trained AudioLDM 2-Music model (Liu et al., 2023c). To ensure a thorough and multidimensional assessment, we measure performance using a suite of metrics that includes CLAP, LPAPS, Structure Distance, and FAD and conduct on an NVIDIA 4090. The computation of these metrics is facilitated by the CLAP models provided within the AudioLDM_eval package, which is publicly available at https://github.com/haoheliu/audioldm_eval.
@@ C.1 Metrics
Objective Metrics There are details about four metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) CLAP Score (Elizalde et al., 2023): This criterion evaluates the degree to which the output conforms to the specified target prompt. (2) Struture Distance (Ju et al., 2023): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio. (3) LPAPS (Iashin & Rahtu, 2021; Paissan et al., 2023): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), this measure evaluates the consistency of the edited audio with the source audio. (4) FAD (Fréchet Audio Distance) (Kilgour et al., 2018): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals.
Subjective Metrics To directly reflect the quality of the audio generated, we carry out MOS (Mean Opinion Score) tests. These tests involve scoring two aspects: MOS-Q, which assesses the edited quality of the audio, and MOS-P, which measures the content preservation of edited audio.
For assessing editing fidelity, the evaluators were specifically directed to “Does the natural language description align with the audio?” They were provided with both the audio and its corresponding caption. They were then asked to give their subjective rating (MOS-Q) on a 20-100 Likert scale.
To assess essential content preservation, human evaluators were presented with source audio, target audio, source prompt, and target prompt. They were then asked to answer the question, “To what extent does the target audio retain the essential features of the source audio, such as melody, instrumentation, and overall style?” The raters had to select one of the options: “completely,” “mostly,” or “somewhat,” using a 20-100 Likert scale for their response.
Our crowd-sourced subjective evaluation tests were conducted via Amazon Mechanical Turk where participants were paid \$8 hourly.
@@ Appendix D Quantitative Results
Analyses on Different CFG Scale The lack of systematic experiments that determine the optimal combination of guidance scales for achieving the best editing performance, and analysis of how these guidance scales affect the final consequence in both reconstruction and editing, we conduct this experiment to find the best scales.
@@
Inference Time We compare the inference time of our method with baselines, and the results are compiled in Table 7. MEDIC achieves the comparative inference time with generation models and inversion techniques. We will make an attempt to reduce the inference time of zero-shot music editing in our future work.
Method Inference Time
AudioLDM 2 42.5s
MusicGen 83.3s
SDEdit 44.3s
DDIM Inversion 81.6s
MusicMagus 89.0s
DDPM-Friendly 33.3s
MEDIC 92.0s
Table 7: Inference Time across different methods.
@@ Appendix F Limitations
In spite of the remarkable outcome of our method, due to the limitation of the generation model we used, we are incapable of instigating a profound change.
Due to the numerous steps it requires (T=200), the duration of computing distance is quite long. Thus, we will implement a more powerful text-to-music generation model to support better editing, while trying to use a consistency model or flow-matching model to achieve high-quality and fast music generation in future work. We will make an attempt to edit more interesting and complex music tasks in the future.
@@ G.1 Change Instrument
In Figure 5, we show the capability of MEDIC to change the instrument. Here we edit the ground truth music piece with the source prompt “a live recording of ambient acoustic [guitar] music" and editing prompt “a live recording of ambient acoustic [violin] music". The difference in instruments can be observed in the Mel-spectrum.
@@ G.2 Add Instrument
In Figure 6, we show the capability of MEDIC to add more instruments. Here we edit the ground truth music piece with the source prompt “a heavy metal instructional audio with a distortion guitar" and editing prompt “a heavy metal instructional audio with a distortion guitar [and drums]". The appearance of the new instrument can be observed in the Mel-spectrum which presents a drum sound of high frequency.
@@
Figure 6: Editing Type 1 Add Instrument
@@ G.3 Delete Instrument
In Figure 7, we show the capability of MEDIC to delete instruments. Here we edit the ground truth music piece with the source prompt “a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood" and the editing prompt “a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood". The vanishing of the instrument can be observed in the Mel-spectrum.
@@
Figure 7: Editing Type 2 Delete Instrument
@@ G.4 Change Genre
In Figure 8, we show the capability of MEDIC to change the genre of a music piece. Here we edit the ground truth music piece with the source prompt “a recording of a solo electric guitar playing [blues] licks" and the editing prompt “a recording of a solo electric guitar playing [rock] licks". The obvious difference in genre can be observed in the Mel-spectrum.
@@
Figure 8: Editing Type 3 Change Genre
@@ G.5 Change Mood
Mood is an important attribute of music. In Figure 9, we show the capability of MEDIC to change the mood of a music piece. Here we edit the ground truth music piece with the source prompt “a recording of [aggressive] electronic and video game music with synthesizer arrangements" and editing prompt “a recording of [peaceful] electronic and video game music with synthesizer arrangements". The change of mood can be observed in the Mel-spectrum.
@@
Figure 9: Editing Type 4 Change Mood
@@ G.6 Change Rhythm
Rhythm represents the speed of the music. In Figure 10, we show the capability of MEDIC to change the Rhythm of a music piece. Here we edit the ground truth music piece with the source prompt “a [slow] tempo ukelele tuning recording with static" and the editing prompt “a [fast] tempo ukelele tuning recording with static". The change of Rhythm can be observed in the Mel-spectrum. The edited Mel-spectrum is much more intensive.
@@
Figure 10: Editing Type 5 Change Rhythm
@@ G.7 Change Background
In Figure 11, we show the capability of MEDIC to change the background of the instrument of a music piece. Here we edit the ground truth music piece with the source prompt “an amateur ukulele recording with a [medium to uptempo] pace" and editing prompt “an amateur ukulele recording with a [steady and rhythmic] pace". The change of instrument background can be observed in the Mel-spectrum.
@@
Figure 11: Editing Type 6 Change Background
@@ G.8 Change Melody
In Figure 12, we show the capability of MEDIC to change the melody of a music piece. Here we edit the ground truth music piece with the source prompt “this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo" and editing prompt “this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo". The change of Melody can be observed in the Mel-spectrum.
@@
Figure 12: Editing Type 7 Change Melody
@@ G.9 Extract Instrument
In Figure 13, we show the capability of MEDIC to extract one certain instrument of a music piece. Here we edit the ground truth music piece with the source prompt “a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]" and editing prompt “a reggae rhythm recording with bongos". The change of instruments can be observed in the Mel-spectrum.
@@
Figure 13: Editing Type 8 Extract Instrument
@@
In the processing of the data and models involved in this study, we fully considered the potential risks. We ensure that all data sources are rigorously screened and vetted, and the model we used is absolutely trained from the safe dataset to minimize the security risks of being misused.